A 3.25-kg block is sent up a ramp inclined at an a
Answer #1

mg*d* sin theta = 1/2*mv^2 - mu_k*m*g*costheta *d

mu_k*m*g*costheta *d = 1/2*mv^2 - mg*d*sintheta

mu_k*g*costheta *d = 1/2 * v^2 - gd sintheta

0.460 * 9.8 * cos37.5 * d = 1/2 *(15)^2 - 9.8 * sin37.5 * d

3.576 d = 112.5 - 5.966d

d = 11.75 m